OpenStudy (anonymous):

Use the Law of Sines to find the length of FE to the nearest tenth of a foot

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\frac{EF}{\sin(39)}=\frac{7}{\sin(34)}\] is a start

OpenStudy (anonymous):

that makes \[EF=\frac{7\sin(39)}{\sin(34)}\] and a calculator to finish

OpenStudy (anonymous):

7.9